OpenStudy (anonymous):

What is the equation for a circle with a center(8,-9) and passes through(21,22)

OpenStudy (aravindg):

first find the radius of the circle

OpenStudy (aravindg):

Use distance formula

OpenStudy (anonymous):

how do i find the radius?

OpenStudy (aravindg):

Distance formula

OpenStudy (aravindg):

you have 2 points : (8,-9) ,(21,22)

OpenStudy (anonymous):

is there a formula to get the radius without graphing it

OpenStudy (aravindg):

ya radius = \[\sqrt{(21-8)^2+(22+9)^2}\]

OpenStudy (whpalmer4):

notice that AravindG has done \(\sqrt{(21-8)^2 + (22-(-9))^2} = \sqrt{(21-8)^2+(22+9)^2}\)

OpenStudy (aravindg):

gt it ?

OpenStudy (whpalmer4):

Once you have the radius, do you know the formula for a circle with center at \((h,k)\) and radius \(r\)?

OpenStudy (anonymous):

i dont know it

OpenStudy (whpalmer4):

Okay, I can tell you the formula, or I can walk you through figuring it out, your choice...

OpenStudy (whpalmer4):

you'll probably remember it better if we go through figuring it out, but it will take a little while.

OpenStudy (whpalmer4):

Well, here's the formula: \[(x-h)^2 + (y-k)^2 = r^2\] Remember that a circle is just the set of points the same distance from its center. If this formula looks suspiciously like the Pythagorean theorem to you, give yourself a gold star.  Subtracting \(h\) and \(k\) before squaring just move the center of the circle around.